OpenStudy (anonymous):

an office has 60 employees 48 have cellular phones what portion of employees so not have cellular phone... helpppppppppp

OpenStudy (anonymous):

20%

OpenStudy (anonymous):

it can be a whole number or fraction only

OpenStudy (anonymous):

60-48=?

OpenStudy (anonymous):

portion of employees so not have cellular phone=60-48=?

OpenStudy (anonymous):

i sis that last time and got it wrong

OpenStudy (anonymous):

did

OpenStudy (anonymous):

60-48= 12 So 12 employees do not have cellular phone.

OpenStudy (anonymous):

whats 20 percent as a fraction

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

((12/60)* 100 = (1/5)*100 = 20% Hence , 20% employees do not have cellular phone.

OpenStudy (anonymous):

1/5